this is from a worksheet covers 10.7 10.8 chp.11
if f(x)= 4/x-2 and g(x)=3/4x, find (f o g )(x), (g o f)(x), and the domain

Answer by stanbon(75887)

I'll do one; you do the other:
f(g(x))= f((3/4)x)=4/[(3/4)x - 2]
Domain?
(3/4)x-2 cannot be zero.
So, if (3/4)x-2=0
(3/4)x=2
x=8/3
So, the Domain is all real numbers except x=8/3
Cheers,
Stan H.